OpenStudy (anonymous):

1. Find the general solution (or as close as you can come to it) for the following differential equations, using separation of variables. (a) dy/dt = 2y − 1 (b) dy/dt = t^2y^3

OpenStudy (amistre64):

separation is just splitting up same variables on same sides

OpenStudy (amistre64):

\[\frac{dy}{dt}=2y-1\] \[dy=(2y-1)dt\] \[(2y-1)^{-1}dy=dt\] \[\int((2y-1)^{-1}dy=dt)\]

OpenStudy (anonymous):

so the asnwer would be \[y=\sqrt[3]{pmAe^(t^3/3)}\]

OpenStudy (anonymous):

that e^(t^3/3)

OpenStudy (amistre64):

\[\int((2y-1)^{-1}dy=dt)\] \[\frac{1}{2}ln|2y-1|=t+C\] \[ln|2y-1|=2t+C\] \[2y-1=\pm C*exp(2t)\] \[y=\pm \frac{Ce^{2t}+1}{2}\] is what I get